OpenStudy (anonymous):

Rationalize the denominator of 9/3 + 2i

OpenStudy (anonymous):

mult top and bottom by the complex conjugate (3-2i)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do I do next

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

27 + 18i/5

OpenStudy (anonymous):

over 5?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

9/(3+2i) =9*(3-2i)/(3+2i)*(3-2i) =9*(3-2i)/(9+4) =9*(3-2i)/11

OpenStudy (anonymous):

\[\frac{ 9 }{ 3 + 2i }\times \frac{ 3 - 2i }{ 3 - 2i }=\frac{ 27 - 18i }{ 9 + 4 }\]

OpenStudy (anonymous):

I see I made a mistake

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no worries, keep that trick in your bag... it'll come in handy on occasion.